Circular No. 8963

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 O2 (McNAUGHT) R. H. McNaught reports his discovery of a comet on CCD images ob-

tained with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below), the comet showing a very condensed 10′′ coma and a tail 1′.0 long in p.a. 240^{o}. Following posting on the ‘NEOCP’ webpage, A. Boattini reports that CCD images taken with the Mt. Lemmon 1.5-m reflector on July 29.4 UT show a 10′′–12′′ coma and a tail almost 2′ long in p.a. 245^{o}. Co-added 60-s CCD images taken on July 29.1 by G. Muler (Lanzarate, Spain, 0.30-m reflector) and measured by G. Sostero show a sharp coma ∼ 12′′ in diameter and a tail nearly 30′′ long toward p.a. 255^{o}. Stacked 60-s CCD images taken on July 29.2 by R. Holmes (Charleston, IL, U.S.A., 0.81-m reflector; measured by L. Buzzi, S. Foglia, and himself) show a narrow tail 20′′ long in p.a. 245^{o}; additional stacked images from Aug. 1.2 show a 30′′ tail in p.a. 190^{o}. J. C. Pelle reports that CCD images obtained by N. Teamo (Punaauia, French Polynesia, 0.35-m telescope) on July 30.35–30.37 show a coma with diameter ∼ 23′′ and a tail > 3′ long in p.a. 240^{o}. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector) writes that images taken on July 29.97 in poor conditions show a coma diameter of 6′′ and a tail 15′′ long in p.a. 250^{o}, while in better conditions on July 30.97, the comet showed an obvious, poorly-condensed coma of diameter 9′′ and a 35′′ tail in p.a. 240^{o}. L. Buzzi (Varese, Italy, 0.60-m f/4.6 reflector) reports a 10′′ coma and a 10′′–15′′ tail in p.a. 230^{o} on stacked images obtained on July 31.9. J. M. Aymami (Tiana, Spain, 0.25m f/4.3 Schmidt-Cassegrain reflector) notes that his stacked CCD images from July 31.9 show a tail ≈ 10′′ long in p.a. 220^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
July 28.69229  21^{h}30^{m}58.^{s}41  −17^{o}00′49′′.6  17.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-P04. This comet is likely of short orbital period, but the orbit is still very indeterminate; additional astrometry is urgently needed.

T = 2009 Nov. 13.429 TT  ω = 118.^{o}098
Ω = 320.457 }2000.0
q = 1.33235 AU  i = 28.676
2008 August 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green